# Circuit-tuning: A Mechanistic Approach for Identifying Parameter Redundancy and Fine-tuning Neural Networks

### v1

@@ 5.1.6 Improvement on Attribution Patching
Acc@1 (%) $\uparrow$ F1 (%) $\uparrow$ Prejudice Risk $\downarrow$ Exact Match / F1 (%) $\uparrow$ Avg param ratio
Llama-3.2-1B-it 40.71 20.35 0.555 39.68, 43.58 /
Llama-3.2-1B-it-full-tuning 46.47 26.89 0.533 36.73, 41.51 1.00
Llama-3.2-1B-it-lora 44.58 22.51 0.530 34.30, 39.11 1.79e-2
Llama-3.2-1B-it-circuit-tuning 45.56 27.06 0.312 41.78, 45.45 7.65e-2
Llama-3.2-3B-it 70.36 42.71 0.641 54.12, 47.98 /
Llama-3.2-3B-it-full-tuning 75.44 47.32 0.632 55.63, 50.10 1.00
Llama-3.2-3B-it-lora 73.54 46.27 0.638 54.73, 49.44 1.49e-2
Llama-3.2-3B-it-circuit-tuning 74.35 47.59 0.417 56.58, 50.93 8.57e-2
Llama-3.1-8B-it 76.19 46.41 0.651 58.92, 52.83 /
Llama-3.1-8B-it-full-tuning 83.76 49.53 0.640 59.60, 54.23 1.00
Llama-3.1-8B-it-lora 80.21 47.64 0.643 58.97, 53.10 1.03e-2
Llama-3.1-8B-it-circuit-tuning 82.97 50.54 0.420 60.04, 55.00 9.37e-2
During the analyses, we notice that some heads fail to appear in the circuit. We assume that the attribution score in EAP fails to measure the importance of a node. This is because an important node $n_{a}$ may be connected with many nodes through edges $\mathcal{E}_{a}$ with low contributions, while another less important node $n_{b}$ may connect with only one node through an edge that is stronger than anyone in $\mathcal{E}_{a}$. As a result, the edges in $\mathcal{E}_{a}$ are discarded, and thus the importance of $n_{a}$ is underestimated. To solve this problem, we revise the contribution $c(e)$ of an edge $e:n_{i}^{u}\rightarrow n_{j}^{d}$ as:
$c(e)^{\prime}=c(e)\cdot\sum_{k=1}^{N_{down}^{i}}c(n_{i}^{u}\rightarrow n_{k}^{d})\cdot\sum_{k=1}^{N_{up}^{j}}c(n_{k}^{u}\rightarrow n_{j}^{d})$
The revision takes into account all the edges connected to the upstream node and the downstream node. The results in Figure 1(c) show that our improvement is effective without introducing too much computation. Details of the results can be found in Section C.4.5.
5.2 Application to Complex Tasks
5.2.1 Task Description
We prepare two types of tasks based on whether reasoning is involved, each of which contains two specific tasks.
For reasoning-based tasks, chain-of-thought (CoT) style responses are involved to solve a problem. We prepare a mathematical task and a logical reasoning task since (Sprague et al., 2024) demonstrated that CoT is effective only on tasks requiring mathematical, logical, or algorithmic reasoning.
@@ 5.2.1 Task Description
For reasoning-free tasks, only the final answer or a signal token is required. We prepare a gender de-biasing task which requires the language model to develop an unbiased perspective on genders, and a reading comprehension task which requires only the keywords as the answer.
As for the metric $\mathcal{L}_{m}$ for circuit discovery, for gender de-biasing, we use the logit difference between male attribution words like $he/his$ and female attribution words like $she/her$. For other tasks, we simply set $\mathcal{L}_{m}$ as identical to the negative log probability loss for language modeling since the behaviors for completing the task are hidden inside multiple tokens and cannot be separated out easily.
As for the loss $\mathcal{L}$, we selectively add $\beta\cdot|\mathcal{L}_{m}|$ as a regularization term for gender de-biasing to improve performance.
Details of task settings can be found in Section D.1.
5.2.2 Dataset and Evaluation Metrics
Reasoning-based Tasks We use GSM8K (Cobbe et al., 2021) with zero-shot accuracy and Contexthub (Hua et al., 2024) with F1 score as datasets and metrics for mathematics and logical reasoning respectively.
Reasoning-free Tasks For the gender-debiasing task, we use BUG (Levy et al., 2021) for training and WinoBias (Zhao et al., 2018) for evaluation. We use the prejudice risk proposed in (Liu et al., 2024) as the evaluation metric. For the reading comprehension task, we use SQuAD 2.0 (Rajpurkar et al., 2018) with exact match and F1 score.
General Capabilities To check if other capabilities are preserved after training, we evaluate on benchmarks involving general abilities as well as reasoning, coding, and multilingual abilities. For details, please refer to Section D.3.
5.2.3 Implementation
We apply our method to llama-3.2-1B / 3B and llama-3.1-8B (Dubey et al., 2024), and compare it with full fine-tuning and LoRA. The graph settings are the same as before in Section 5.1.2, except that each MLP layer is split into 64-dimensional heads. The details are shown in Section D.2.
5.2.4 Main Results
Figure 3: Comparison of different fine-tuning methods
Results of our method on complex tasks in Table 1 demonstrate that our method is capable of tasks in various scenarios. Note that with a regularization term as guidance, the de-biasing result is clearly better than others, which is further shown in Section D.4. In general, circuit-tuning is better than LoRA, and even surpasses full fine-tuning in many tasks with only a small portion of tunable parameters.
Evaluations on general capabilities are shown in Figure 3. The results are averaged over tasks. It can be seen that our method keeps comparable performance with full fine-tuning while preserving general capabilities better than others.
@@ 6 Conclusion
In conclusion, we propose the theory of node redundancy from the mechanistic view as an extension of the intrinsic dimension. With the guidance of the theory, we propose the algorithm of circuit-tuning, which not only shows great potential in fine-tuning but also serves as a useful tool for analyzing the training dynamics inside a model.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. Since our work concerns AI alignment and involves modifying a model’s gender stereotype, we expect our work to be used under the guidance of human values.
Appendix A The Derivation of Edge Contribution
According to the definition of attribution score in (Syed et al., 2023) and the definition of node redundancy in Section 3.3, we can define the contribution of an edge $e:n_{1}\rightarrow n_{2}$ with $n_{1}$ as the upstream node and $n_{2}$ as the downstream node. We use the indirect effect $IE$ to measure the change in the output caused by the patching of the edge. Thus given a dataset $\mathcal{X}$ for patching, the contribution of edge $e$ can be expressed as follows:
@@ Appendix A The Derivation of Edge Contribution
$=\mathbb{E}_{x_{i}\sim\mathcal{X}}\big{[}|f(e;M(x_{i}),M(x_{i}^{\prime}))|\big{]}$
$=\mathbb{E}_{x_{i}\sim\mathcal{X}}\big{[}|IE(e;x_{i})|\big{]}$
Note that the contribution of edge $e$ is directly reflected in the change of the final output (the logit of the language model, etc), which is the indirect effect caused by the change of the value in the downstream node $n_{2}$, while the change of the node $n_{2}$ is directly caused by the change of the upstream node $n_{1}$. The difference between the direct effect and the indirect effect is that the former keeps all the other nodes that could influence $n_{2}$ unchanged and only studies the influence from $n_{1}$ to $n_{2}$, while the latter allows all the changes in nodes between $n_{2}$ and the logit. For more refined definitions for these two concepts, please refer to (Pearl, 2001).
@@
$=$ $\mathcal{L}_{m}\big{[}M\big{(}x\,|\,do\big{(}n_{2}\leftarrow n_{2}(x^{\prime})\big{)}\big{)}\big{]}-\mathcal{L}_{m}\big{[}M(x)\big{]}$
@@
$n_{2}^{n_{1}}(x^{\prime})=n_{2}^{n_{1}}\big{(}x\,|\,do\big{(}n_{1}\leftarrow n_{1}(x^{\prime})\big{)}\big{)}$
@@
$IE(e;x)\approx$ $\mathcal{L}_{m}[M(x)]+\big{[}n_{2}(x^{\prime})-n_{2}(x)\big{]}^{\top}\nabla_{n_{2}}\mathcal{L}_{m}[M(x)]\big{|}_{n_{2}(x)}-\mathcal{L}_{m}[M(x)]$
$=$ $\big{[}n_{2}(x^{\prime})-n_{2}(x)\big{]}^{\top}\nabla_{n_{2}}\mathcal{L}_{m}[M(x)]\big{|}_{n_{2}(x)}$
@@
To further simplify Appendix A, we apply another Taylor expansion at $n_{1}=n_{1}(x)$ to $n_{2}^{n_{1}}$. Then we have

### v2

@@ 5.2 Application to Complex Tasks
Acc@1 (%) F1 (%) Prejudice Risk $\downarrow$ Exact Match(%) Avg param ratio
Llama-3.2-1B-it 40.71 20.35 0.555 39.68 /
Llama-3.2-1B-it-full-tuning 46.47 26.89 0.533 36.73 1.00
Llama-3.2-1B-it-lora 44.58 22.51 0.530 34.30 1.79e-2
Llama-3.2-1B-it-circuit-tuning 45.56 27.06 0.312 41.78 7.65e-2
Llama-3.2-3B-it 70.36 42.71 0.641 54.12 /
Llama-3.2-3B-it-full-tuning 75.44 47.32 0.632 55.63 1.00
Llama-3.2-3B-it-lora 73.54 46.27 0.638 54.73 1.49e-2
Llama-3.2-3B-it-circuit-tuning 74.35 47.59 0.417 56.58 8.57e-2
Llama-3.1-8B-it 76.19 46.41 0.651 58.92 /
Llama-3.1-8B-it-full-tuning 83.76 49.53 0.640 59.60 1.00
Llama-3.1-8B-it-lora 80.21 47.64 0.643 58.97 1.03e-2
Llama-3.1-8B-it-circuit-tuning 82.97 50.54 0.420 60.04 9.37e-2
Task descriptions
@@
We prepare two types of tasks based on whether reasoning is involved. For reasoning-based tasks, we focus on mathematics and logical reasoning with chain-of-thought (CoT) included. For reasoning-free tasks, we prepare a gender de-biasing task that requires the model to develop an unbiased perspective on genders, and a reading comprehension task that requires only the keywords as the answer. Task settings and evaluation metrics are detailed in Appendix G.1.
Figure 5: Comparison in gender debiasing results.
Figure 6: Evaluation on general/other capabilities.
Implementation
We apply our method to llama-3.2-1B/3B and llama-3.1-8B [11]. In addition, we compare it with full fine-tuning and LoRA. We set them as baselines for simple comparisons only to verify that our method is also able to locate the key parameters precisely, as discussed in Section 4.2.
For circuit-tuning, the graph settings of our method are the same as before in Section 5.1.2, except that each MLP layer is split into 64-dimensional heads. For gender de-biasing, We selectively add $\beta\cdot|\mathcal{L}_{m}|$ that acts as a regularization term in the loss $\mathcal{L}$ for explicit guidance and intervention during training (Appendix G.1.2). All other details like the choice of the patching metric $\mathcal{L}_{m}$, the number of edges $N$, and the hyper-parameters for circuit-tuning and baselines are detailed in Appendix G.2.
Main results Table 2 shows the results of each method, demonstrating that our method is capable of various scenarios. Hence it is possible to study the learning dynamics as in Section 5.1 on various scenarios. Amazingly, the task performance brought by circuit-tuning is better than LoRA, and even surpasses full fine-tuning in many tasks. Note that the computation of circuit-tuning depends on node granularity and is for reference only since our method is not dedicated to efficient fine-tuning.
Besides, with an interpretable intervention, the de-biasing result is clearly better than others, which is shown in Figure 6 and discussed in Appendix G.4. This suggests that circuit-tuning can be applied flexibly with task-specific conditions to achieve controllable and steerable training.
To check if other capabilities are affected during training, we evaluate on benchmarks involving general capabilities as well as reasoning, coding, and multilingual abilities. Results are shown in Figure 6 and detailed in Appendix G.3, indicating that circuit-tuning is more targeted at the key components in a model, ensuring its reliability and simplicity in interpretability research.
@@ 6 Conclusion
We extend the concept of intrinsic dimensionality to the node level, and describe the learning process inside a model from the mechanistic view. Further, we propose circuit-tuning as a promising tool for interpretable fine-tuning. We analyze the mechanism behind learning and provide new insights for the learning process. Limitations and future discussions of our work are shown in Appendix A.
Appendix A Limitations and future discussions
In this section, we report some limitations about our work, and discuss some potential research directions based on our method.
  1. 1.
    We do not test our method with different granularities. The node in a computational graph could be a neuron, a group of neurons, the activation of an attention head or even a layer. Also, it could be a latent in the representation of a sparse autoencoder. While in our experiment, we treat the activation of each attention / MLP head as a node for convenience. We believe it is possible to try other granularities.
  2. 2.
    We only focus on attention heads as most people do. In fact, we believe that a lot of treasures are hidden inside the MLP layers. Recently Ameisen et al. [2], Lindsey et al. [32] applied transcoders to MLP layers and discovered the circuits inside a model. Thus we look forward to seeing circuit-tuning combined with other model components besides attention.
  3. 3.
    We believe that the experiment on complex tasks could be further studied. For example, we could analyze the learning dynamics during the fine-tuning process for COT-based tasks, or COT-free tasks like the gender de-biasing task. However, due to the complexity of these tasks and limited space, we think it is a better idea to study them in a separate work and present them in a single paper. Thus we leave this part for further research.
  4. 4.
    The infra of our method could be improved. We believe it is possible to apply our method to larger models, while it requires stronger frameworks. We leave it for furture development.
Appendix B Social impact
This paper presents work whose goal is to advance the field of Machine Learning. Since our work concerns AI alignment and involves modifying a model’s gender stereotype in Section 5.2, we expect our work to be used under the guidance of human values.
Appendix C Transformer Architecture
A single input of Transformer is $x_{0}\in\mathbb{R}^{T}$, where $T$ is the length of the sequence. The input is firstly embedded into a vector $x\in\mathbb{R}^{D\times T}$ via the embedding matrix $W_{E}\in\mathbb{R}^{D\times V}$, where $D$ is the model dimension. Then $x$ will go through $l$ layers of Transformer blocks for various processings. From the view of [13], we can think of the residual stream as a communication channel that simply receives the output of the self-attention and MLP operations. Each operation reads information from the residual stream and writes the processed information into it. Thus the residual stream is actually the linear sum of various transformations of $x$ together with the original input $x$.
For the MLP operation in each layer, the input $x$ is projected into $x_{in}^{i}\in\mathbb{R}^{D_{mlp}\times T}$ via $W_{in}\in\mathbb{R}^{D_{mlp}\times D}$, and projected back to $MLP_{i}(x)\in\mathbb{R}^{D\times T}$ via $W_{out}\in\mathbb{R}^{D\times D_{mlp}}$. In the Llama architecture [59], the input $x$ is firstly projected into $x_{pre}^{i}\in\mathbb{R}^{D_{mlp}\times T}$ via $W_{gate}^{i}\in\mathbb{R}^{D_{mlp}\times D}$ and is applied with an activation layer, then a dot product is performed between the activations and $x_{in}^{i}\in\mathbb{R}^{D_{mlp}\times T}$ which is the input projected by $W_{in}\in\mathbb{R}^{D_{mlp}\times D}$. Note that we can also split the MLP into MLP heads, which is done on Llama series models in the complex tasks in our experiments. For details, please refer to Appendix G.2.
Appendix D The Derivation of Edge Contribution
Similar to the definition of attribution score in [56], we can define the contribution of an edge $e:n_{1}\rightarrow n_{2}$ with $n_{1}$ as the upstream node and $n_{2}$ as the downstream node. We use the indirect effect $IE$ to measure the change in the output caused by the patching of the edge. Thus given a dataset $\mathcal{X}$ for patching, the contribution of edge $e$ can be expressed as follows:
@@ Appendix D The Derivation of Edge Contribution
$=\mathbb{E}_{x_{i}\sim\mathcal{X}}\big{[}|IE(e;x_{i})|\big{]}$ (4)
Note that the contribution of edge $e$ is directly reflected in the change of the final output (the logit of the language model, etc), which is the indirect effect caused by the change of the value in the downstream node $n_{2}$, while the change of the node $n_{2}$ is directly caused by the change of the upstream node $n_{1}$. The difference between the direct effect and the indirect effect is that the former keeps all the other nodes that could influence $n_{2}$ unchanged and only studies the influence from $n_{1}$ to $n_{2}$, while the latter allows all the changes in nodes between $n_{2}$ and the logit. For more refined definitions for these two concepts, please refer to [44].
@@
$=$ $\mathcal{L}_{m}\big{[}M\big{(}x\,|\,do\big{(}n_{2}\leftarrow n_{2}(x^{\prime})\big{)}\big{)}\big{]}-\mathcal{L}_{m}\big{[}M(x)\big{]}$ (5)
@@
$n_{2}^{n_{1}}(x^{\prime})=n_{2}^{n_{1}}\big{(}x\,|\,do\big{(}n_{1}\leftarrow n_{1}(x^{\prime})\big{)}\big{)}$ (7)
@@
$IE(e;x)\approx$ $\mathcal{L}_{m}[M(x)]+\big{[}n_{2}(x^{\prime})-n_{2}(x)\big{]}^{\top}\nabla_{n_{2}}\mathcal{L}_{m}[M(x)]\big{|}_{n_{2}(x)}-\mathcal{L}_{m}[M(x)]$ (8)
$=$ $\big{[}n_{2}(x^{\prime})-n_{2}(x)\big{]}^{\top}\nabla_{n_{2}}\mathcal{L}_{m}[M(x)]\big{|}_{n_{2}(x)}$ (9)
@@
To further simplify Equation D, we apply another Taylor expansion at $n_{1}=n_{1}(x)$ to $n_{2}^{n_{1}}$. Then we have